# Shortcut Bias Mitigation via Ensemble Diversity Using Diffusion Probabilistic Models

### 23 Nov 2023 v1

@@ 5 Discussion and Conclusion
Summary. Shortcut learning is a phenomenon undermining the performance and utility of deep learning models, where easy-to-learn cues are preferentially learned for prediction, regardless of their relevance to the downstream task. We investigate the potential of DPMs to aid in the mitigation of this phenomenon by training a diverse ensemble through disagreement on synthetic counterfactuals. We show several important findings.
First, we show that DPMs can generate novel feature combinations even when trained on images displaying correlated cues, and that this phenomenon is more prevalent at particular training intervals. One significant implication of these findings is the potential of DPMs for feature disentanglement and ood sample generation, which may have interesting repercussions in several important domains including data augmentation and ood generalization.
@@
Second, we show that diffusion-guided diversification leads models to avert attention from shortcut cues, and that diffusion counterfactuals can lead to comparable ensemble diversity without the need for expensive ood data collection.
Limitations and Future Directions. The results, and limitations, of this study open several avenues for future research. Firstly, it would be useful to understand the mechanism behind the ability of DPMs to generalize beyond the observed feature combinations, even when trained on images displaying fully correlated input features.
In our study, we find that there is a delicate balance between the level of diffusion fidelity and the effectiveness of ensemble diversification. In particular, we find it useful for a generative model to learn the manifold of the distribution under scrutiny, without (over)fitting the intricate nuances of the training data. Another interesting direction would be to delve deeper into the interplay between the fidelity of DPM-generated samples and model diversification.
@@
A limiting factor in the disagreement objective to achieve shortcut mitigation lies in its impact on average ensemble performance. In prior work, this was partly overcome by careful model selection, but new avenues should be explored to maintain iid performance while achieving diverse extrapolation trends via disagreement.
@@
Conclusion. This work presents a step forward in addressing the challenge of mitigating shortcut learning in DNNs. By leveraging the unique capabilities of DPMs for ensemble diversification we provide a practical method that achieves shortcut mitigation in two visual tasks, with only negligible performance loss compared to methods requiring expensive ood data collection.
Acknowledgements
Luca Scimeca acknowledges funding from Recursion Pharmaceuticals. Damien Teney is partially supported by an Amazon Research Award.
Supplementary Material
6 Supplementary Results
6.1 Calibration
@@
We perform a hyper-parameter search to find the values of $\gamma$ for each diversification objective. We perform the search via the same methodology used in the ood main experiments, i.e. by training an ensemble of 100 ResNet-18 models on the fully correlated diagonal datasets, while diversifying a small subset of 30\$ of the original training data, randomly sampled from the de-correlated left-out set. We consider $\gamma$ values ranging from $31e-$ 3 to $11e$ 1, and monitor both the validation ensemble accuracy as well as the predictive diversity on a separate de-correlated set of validation data. Figure S1 shows the performance of each metric for different values of $\gamma$. We select the values reported in Table TABLE to be at the intersection of the accuracy and diversification trends for each model.
@@
Figure S1: Hyperparameter search on the $\gamma$ diversification intensity for each diversification objective. For the main experiments, we select $\gamma$s at the intersection of the accuracy-diversity trends by each objective.
6.2 On the influence of DPM Fidelity to diversification
We perform experiments whereby an ensemble of 100 ResNet-18 models is trained separately with respect to all diversification objectives considered. Figure S2 shows the diversification results as a function of the fidelity of the DPM used to sample the diversification set. We observe the diversification level obtained is highly dependent on the Diffusion Fidelity level, and that appropriate early stopping procedures must be followed to achieve increased ensemble prediction diversity.
@@ 6.2 On the influence of DPM Fidelity to diversification
6.3 Diversification Leads to Ensemble Models Attending to Different Cues
Figure S3 illustrates a feature-centric description of 10 ensemble models trained with a diversification objective on OOD data (a) and Diffusion generated counterfactuals (b). Within each results are shown for ColorDSprites (upper panel) and UTKFace (lower panel). The variation across models is evident: several models substantially reduce their dependency on the leading cues of the respective datasets (black edges), diverging considerably from the almost identical configurations present in the baseline ensemble (red edges).
@@ 6.3 Diversification Leads to Ensemble Models Attending to Different Cues
6.4 On the Influence of an Increased Number of OOD Samples for Ensemble Disagreement
As per the original objective, with DPM sampling we aim to circumvent the diversification dependency on Out-Of-Distribution data, which is often not readily accessible and can be costly to procure. We test this dependency further and assess the quality of the diversification results when matching the number of OOD data used for diversification to the original training data for the ensemble. We report in Table S1 our findings. We observe the quality of the disagreement on ColorDSprites to only marginally benefit from additional disagreement samples, with approximately $10$ to $15\%$ of the models to avert their attention from the shortcut cue color. On the other hand, we observe a strong improvement in the diversification for UTKFace, mainly registered via the cross objective, where approximately $39\%$ of the models avert their attention from the ethnicity shortcut, as opposed to the original $1\%$ to $20\%$ in our previous experiments, while maintaining high predictive performance on the validation set. We speculate this gain to be due to the higher complexity of the features within the data, which may require additional specimens for appropriate diversification.

### 05 Mar 2024 v2

@@ 5 Discussion and Conclusion
Shortcut learning is a phenomenon undermining the performance and utility of deep learning models, where easy-to-learn cues are preferentially learned for prediction, regardless of their relevance to the downstream task. We investigate the potential of DPMs to aid in the mitigation of this phenomenon by training a diverse ensemble through disagreement on synthetic counterfactuals
DPM training and ood generation
Summary. We show that DPMs can generate novel feature combinations even when trained on images displaying correlated cues. We observe this phenomenon as a function of the diffusion training epochs, suggesting early tendencies of DPMs to learn the manifold of the distribution under scrutiny, without (over)fitting the intricate nuances of the training data.
Implications and future directions. Although beyond the scope of this work, it would be worthwhile to understand the mechanism behind the ability of DPMs to generalize beyond the observed feature combinations even under feature correlation. An implication of this finding is the potential of early-stopping mechanisms to enforce these particular generative capabilities. Another significant implication is the potential of DPMs for feature disentanglement and ood sample generation, which may have interesting repercussions in several important domains including data augmentation and ood generalization.
Ensemble Diversification
@@
Summary. We show that diffusion-guided diversification leads models to avert attention from shortcut cues, and that diffusion counterfactuals can lead to comparable ensemble diversity without the need for expensive ood data collection. In our study, we find that there is a delicate balance between the level of diffusion fidelity and the effectiveness of ensemble diversification. In particular, we find that ensemble diversity and validation iid performance can be used as a proxy for the identification of the DPM originative stage.
@@
Implications and future directions. We believe it worthwhile to delve deeper into the interplay between the fidelity of DPM-generated samples and model diversification. A limiting factor in the disagreement objective to achieve shortcut mitigation lies in its impact on average ensemble performance. In prior work, this was partly overcome by careful model selection, but new avenues should be explored to maintain iid performance while achieving diverse extrapolation trends via disagreement. Furthermore, although our findings showcase an important phenomenon, and its applicative utility for diverse ensembles, it would also be valuable to extend this work to additional datasets and explore its implications in data modalities beyond vision.
@@
Conclusion. This work presents a step forward in addressing the challenge shortcut learning in DNNs. By leveraging the unique capabilities of DPMs for ensemble diversification we provide a practical method that achieves shortcut mitigation in two visual tasks, with only negligible performance loss compared to methods requiring expensive ood data collection.
S1 Supplementary Methods
S1.1 Operationalizing WCST-ML Across Datasets
S2 Supplementary Results
S2.1 Calibration
@@
We perform a hyper-parameter search to find the values of $\gamma$ for each diversification objective. We perform the search via the same methodology used in the ood main experiments, i.e. by training an ensemble of 100 ResNet-18 models on the fully correlated diagonal datasets, while diversifying a small subset of 30% of the original training data, randomly sampled from the de-correlated left-out set. We consider $\gamma$ values ranging from $31e-$ 3 to $11e$ 1, and monitor both the validation ensemble accuracy as well as the predictive diversity on a separate de-correlated set of validation data. Figure S1 shows the performance of each metric for different values of $\gamma$. We select the values reported in Table S1 to be at the intersection of the accuracy and diversification trends for each model.
@@
Figure S1: Hyper-parameter search on the disagreement intensity ($\gamma$) for each diversification objective. For the main experiments, we select the values of $\gamma$s at the intersection of the accuracy-diversity trends by each objective (Table S1).
S2.2 On the Influence of DPM Fidelity to Diversification
We perform experiments whereby an ensemble of 100 ResNet-18 models is trained separately with respect to all diversification objectives considered. Figure S4 shows the diversification results as a function of the fidelity of the DPM used to sample the diversification set. We observe the diversification level obtained is highly dependent on the diffusion fidelity level, and that appropriate early stopping procedures are necessary to achieve increased ensemble prediction diversity. while the trends vary mildly across different disagreement objectives, we find the diversification maximized in ColorDSprites and UTKFace coherently with the analysis in Section 4.1, where improved diversification is achieved around the originative interval. For ColorDSprites, this is realized in around 20 DPM training epochs, while for UTKFace, it is realized in around 800 DPM training epochs.
Importantly, our results suggest how ensemble diversification metrics can be a viable proxy for appropriate originative DPM training. In Fig. S3, we observe the min-maxed change in accuracy and change in diversity by the ensembles with respect to baseline training. In the figure, we observe similar trends across all diversification methods, whereby the originative interval (highlighted in gray) identified by the highest changes in diversity and least drop in accuracy, highly overlaps with our previous supervised findings (Fig. 3). Our results align with our previous observations, where the highest number of ood-generated samples to lie within the DPM training intervals achieve the highest change in diversity while maintaining good classification performance (Fig. 6.)
Figure S2: ColorDSprites
Figure S3: Min-Maxed change in accuracy and diversity by ensembles trained with diffusion-augmented samples, with respect to all considered diversification methods. The originative stage, as qualitatively identified in the experiments (see Fig. 3 and Fig. 4) is shown in gray.
Table S1: Disagreement $\gamma$ used in our experiments.
l2 l2 cross div kl
color 5.0 1.5 0.5 0.5 0.1
face 5.0 2.0 2.0 5.0 0.2
@@
S2.3 Diversification Leads to Ensemble Models Attending to Different Cues
Figure S5 illustrates a feature-centric description of 10 ensemble models trained with a diversification objective on ood data (a) and Diffusion generated counterfactuals (b). Within each results are shown for ColorDSprites (upper panel) and UTKFace (lower panel). The variation across models is evident: several models substantially reduce their dependency on the leading cues of the respective datasets (black edges), diverging considerably from the almost identical configurations present in the baseline ensemble (red edges). For some of the models, the averted attention on the main shortcut cue leads to increased reliance on one of the other observed features (e.g. scale and age for models 7 (above) and 73 (below) in Fig. 4(a), and models 80 (above) and 47 (below) in Fig. 4(b)). In some, however, the averted attention on the main shortcut cues instead leads to a reliance on extracted image features unobserved in our experiments (e.g. Models 28 (above) and 67 (below) in Fig. 4(a), and models 98 (above) and 48 (below) in Fig. 4(b))).
S2.4 On the Influence of an Increased Number of ood Samples for Ensemble Disagreement
As per the original objective, with DPM sampling we aim to circumvent the diversification dependency on Out-Of-Distribution data, which is often not readily accessible and can be costly to procure. We test this dependency further and assess the quality of the diversification results when matching the number of ood data used for diversification to the original training data for the ensemble. We report in Table S2 our findings. We observe the quality of the disagreement on ColorDSprites to only marginally benefit from additional disagreement samples, with approximately $1\%$ to $7\%$ more of the models to avert their attention from the shortcut cue color as compared to the original experiments. On the other hand, we observe a strong improvement in the diversification for UTKFace, mainly registered via the div objective, where $24\%$ of the models averted their attention from the ethnicity shortcut, as opposed to the original $6\%$ in our previous experiments, while maintaining high predictive performance on the validation set. We observe marginal improvements on the other objectives, with approximately $4\%$ to $8\%$ additional models achieving cue aversion. We speculate this gain to be due to the higher complexity of the features within the data, which may require additional specimens for appropriate diversification.